### v2

@@ 3.2 Confidence-Tuning for Deferral
$\mathcal{L}_{\text{incorr}}$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathds{1}\{y_{i,t}\neq\hat{y}_{i,t}\}\,\text{KL}(p_{i,t}\parallel\mathcal{U})$ (5)
where $T$ is the sequence length, and $p_{i,t}$ is the token distribution at position $t$. This ensures fine-grained calibration across the sequence.
Stage 3: Confidence-Based Deferral. After fine-tuning with Gatekeeper, we use calibrated confidence scores from $\mathcal{M}_{S}$ to make deferral decisions. Following the selective prediction framework (El-Yaniv & Wiener, 2010), we define a gating function $g(\mathbf{x})$ and threshold $\tau$ for routing:
$(\mathcal{M}_{S},\mathcal{M}_{L},g)(\mathbf{x})=\begin{cases}\mathcal{M}_{S}(\mathbf{x})&\text{if }g(\mathbf{x})\geq\tau\\ \mathcal{M}_{L}(\mathbf{x})&\text{otherwise.}\end{cases}$ (6)
For classification models, we use max-softmax confidence:
$g_{\text{CL}}(\mathbf{x})=\max_{c}p(y=c\mid\mathbf{x})$ (7)
For sequence models, we use negative predictive entropy averaged across tokens:
$g_{\text{NENT}}(\mathbf{x})=\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{C}p(y_{t}=c\mid\mathbf{x})\log p(y_{t}=c\mid\mathbf{x})$ (8)
In both cases, higher scores indicate higher confidence. By thresholding this signal, we control which inputs $\mathcal{M}_{S}$ handles autonomously and which are escalated to $\mathcal{M}_{L}$—enabling an efficient, robust cascaded inference system.
@@
4 Experiments
In this section, we detail the experiments used to evaluate the effectiveness of Gatekeeper across three distinct model architectures: encoder-only classification models, decoder-only language models, and encoder-decoder vision–language models. Each setup involves a cascade where a smaller model defers uncertain inputs to a larger model.
4.1 Encoder-only Setup (Classification Models)
We begin our evaluation with image classification using encoder-only models. We train a small model $\mathcal{M}_{S}$ and a larger model $\mathcal{M}_{L}$ on CIFAR-10/100 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and TinyImageNet200 (Le & Yang, 2015). For CIFAR, $\mathcal{M}_{L}$ is a ResNet-18 and $\mathcal{M}_{S}$ is a custom CNN. For Food-101 and TinyImageNet, $\mathcal{M}_{L}$ is a ResNet-50 and $\mathcal{M}_{S}$ is a MobileNet V3 Small (Howard et al., 2019) trained by distilling $\mathcal{M}_{L}$.
@@ 4.1 Encoder-only Setup (Classification Models)
Figure 5: Performance trade-off between small model accuracy $\text{acc}(\mathcal{M}_{S})$ and deferral performance $s_{d}$. The baseline model obtained without fine-tuning using Gatekeeper is often the most accurate model over the full data distribution. With the introduction of Gatekeeper we can improve distinguishability of correct/incorrect predictions (left) as well as deferral (right) at the expense of model utility. Successful cascading solutions in practice need to balance both model accuracy and deferral performance.
Evaluation Metrics. We evaluate the effectiveness of Gatekeeper and the resulting deferral function $g(\cdot)$ using three metrics that reflect different aspects of performance (see Figure 3 for an illustrative example):
  1. 1.
    Distributional Overlap of Confidence Scores $s_{o}$: Measures how well $\mathcal{M}_{S}$ separates correct from incorrect predictions based on output confidence. We estimate KDEs over confidence scores for correctly ($\hat{p}_{\text{corr}}$) and incorrectly ($\hat{p}_{\text{incorr}}$) classified samples and define:
@@
    Lower $s_{o}$ indicates better separability ($s_{o}=0$ for perfect, $s_{o}=1$ for complete overlap). This metric is related to AUROC but compares full probability mass rather than prediction ranking.
  2. 2.
    Deferral Performance $s_{d}$: Quantifies how effectively $\mathcal{M}_{S}$ defers uncertain inputs to $\mathcal{M}_{L}$. We compute the joint accuracy $\text{acc}_{\text{real}}(r)$ over varying deferral rates $r$, and compare it against two baselines: random deferral $\text{acc}_{\text{rand}}(r)$ and ideal deferral $\text{acc}_{\text{ideal}}(r)$. The score is the normalized area between real and random deferral:
@@
    A value of $s_{d}=1$ indicates optimal deferral; $s_{d}=0$ implies no improvement over random routing. See Appendix B.3 for details.
  3. 3.
    Small Model Accuracy $\text{acc}(\mathcal{M}_{S})$: Since Gatekeeper prioritizes uncertainty calibration over classification accuracy, $\mathcal{M}_{S}$ may sacrifice performance on the full data distribution to better recognize its errors. We report $\text{acc}(\mathcal{M}_{S})$ to quantify this trade-off and assess utility loss when optimizing for deferral.
Results. Figure 4 shows results for models trained with varying $\alpha$ values, compared against a baseline model and the cascading method of Narasimhan et al. (2022). At low $\alpha$, Gatekeeper effectively separates correct and incorrect confidence scores (lower $s_{o}$), improving deferral performance $s_{d}$. However, this comes with reduced small model accuracy, as $\mathcal{M}_{S}$ increasingly focuses on easy examples and assigns lower confidence to hard ones. As $\alpha$ increases, accuracy improves or stabilizes, but gains in deferral diminish. Notably, the baseline from Narasimhan et al. (2022) maintains model accuracy but requires auxiliary mechanisms to predict expert correctness—such as additional heads or networks—which can increase deployment complexity. In contrast, Gatekeeper is architecture-agnostic and operates purely via confidence calibration.
The accuracy-deferral trade-off is further explored in Figure 5, where we observe: (i) a negative correlation between deferral performance $s_{d}$ and small model accuracy, and (ii) a strong relationship between confidence overlap $s_{o}$ and both deferral and accuracy. The effect mirrors well-known trade-offs in fairness (Dutta et al., 2020; Yaghini et al., 2023) and privacy (Abadi et al., 2016; Rabanser et al., 2023). Crucially, Gatekeeper exposes this trade-off explicitly via a single tunable parameter $\alpha$, allowing practitioners to tailor cascaded systems to their computational or reliability needs.
Figure 7: Performance on VLM classification (left) and captioning tasks (right). Consistent with results in Figures 4 and 6, we see that smaller $\alpha$s lead to improved deferral performance.
4.2 Decoder-only Setup (Language Models)
We next evaluate Gatekeeper on decoder-only language models. Our setup uses Gemma2B as $\mathcal{M}_{S}$ and Gemma7B as $\mathcal{M}_{L}$ (GemmaTeam et al., 2024), forming a scalable cascade for next-token prediction. As in the classification case, the deferral decision is based solely on the output of $\mathcal{M}_{S}$, using predictive entropy to route uncertain tokens to $\mathcal{M}_{L}$.
Both models are first instruction-tuned on the training split of each dataset to ensure strong base performance and adherence to task format. We then fine-tune $\mathcal{M}_{S}$ with Gatekeeper to reduce its confidence on incorrect token predictions. Evaluation is performed on the validation split using the same metrics as in Section 4.1. We use three representative benchmarks: ARC-e/c (Clark et al., 2018), MMLU (Hendrycks et al., 2020), and GSM8K (Cobbe et al., 2021). These span symbolic reasoning, factual recall, and multi-step arithmetic generation, respectively.
Results. Figure 6 shows that Gatekeeper leads to improved deferral performance and better separation of correct/incorrect predictions, particularly at lower $\alpha$. As with classification, higher $\alpha$ values preserve $\mathcal{M}_{S}$ accuracy but yield weaker separation. In addition to the baseline (Gemma2B with entropy-based deferral), we compare against: (i) the token-level cascading technique from Gupta et al. (2024); and (ii) two uncertainty prompting baselines (see Appendix C.2): Reduce Confidence and Answer “N”. These methods follow Kadavath et al. (2022) and aim to elicit uncertainty via prompting. Overall, we find that Gatekeeper consistently outperforms both prompting and token-level deferral baselines in terms of deferral quality, without modifying the model architecture or inference pipeline.
4.3 Encoder-Decoder Setup (Vision-Language Models)
We conclude by evaluating Gatekeeper on encoder-decoder vision-language models. Specifically, we use the PaliGemma family (Steiner et al., 2024), which supports tasks like image captioning, VQA, and descriptive classification. $\mathcal{M}_{S}$ is PaliGemma1B, and $\mathcal{M}_{L}$ is PaliGemma7B. Our cascade runs $\mathcal{M}_{S}$ on all inputs and defers to $\mathcal{M}_{L}$ when the predictive entropy of $\mathcal{M}_{S}$ is low. Following Section 4.2, we first fine-tune both models on the task using standard supervised instruction tuning. We then apply Gatekeeper to $\mathcal{M}_{S}$ alone and evaluate the resulting cascade on two classification datasets (VQAv2 (Goyal et al., 2017), AI2D (Hiippala et al., 2021)) and two captioning datasets (Cococap (Lin et al., 2014), Screen2Words (Wang et al., 2021)), covering both closed-form and generative VL tasks.
Factuality Scoring. For classification tasks, we evaluate accuracy and confidence separation as before. For captioning, however, we assess factual alignment between generated and reference captions using the Gemini LLM (GeminiTeam et al., 2023). The model is prompted with: “Are these captions semantically equivalent?” and outputs “Yes” or “No.” We compute the normalized log-likelihood of each response to obtain a factuality score $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$, reflecting the model’s confidence in semantic agreement. Full details are provided in Appendix C.4.
Uncertainty/Factuality Correlation. Since factuality scores are continuous, we adapt our metrics accordingly. Instead of binary accuracy-based separation, we compute the Pearson correlation $\rho\big{(}g_{\text{NENT}}(\mathbf{x}_{i}),s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})\big{)}$ between negative predictive entropy and factuality. We also generalize our deferral performance metric $s_{d}$ to operate on factuality instead of accuracy.
Results. Figure 7 shows our main results. For classification tasks (left), trends match those observed in Sections 4.1 and 4.2: lower $\alpha$ improves deferral but reduces raw accuracy. For captioning tasks (right), Gatekeeper increases the correlation between confidence and factuality, demonstrating effective deferral in generative settings. Prompting baselines from Section 4.2 could not be evaluated: PaliGemma failed to return valid outputs under prompt modifications, likely due to rigid pretraining (Beyer et al., 2024).

### v3

@@ 3.2 Confidence-Tuning for Deferral
$\mathcal{L}_{\text{corr}}=\frac{1}{N}\sum_{\begin{subarray}{c}i=1\\ t=1\end{subarray}}^{\begin{subarray}{c}N,\;T\end{subarray}}\mathds{1}\{y_{i,t}=\hat{y}_{i,t}\}\text{CE}(p_{i,t}(\mathbf{x}_{i}),y_{i,t})\qquad\mathcal{L}_{\text{incorr}}=\frac{1}{N}\sum_{\begin{subarray}{c}i=1\\ t=1\end{subarray}}^{\begin{subarray}{c}N,\;T\end{subarray}}\mathds{1}\{y_{i,t}\neq\hat{y}_{i,t}\}\text{KL}\left(p_{i,t}(\mathbf{x}_{i})\parallel\mathcal{U}\right).$ (3)
Here, $y_{i,t}$ and $\hat{y}_{i,t}$ denote the true and predicted tokens at position $t$ for sample $i$, $p_{i,t}$ is the predicted token distribution at position $t$ for sample $i$, and $T$ is the sequence length for the token-based model. The token-level loss ensures that correct token predictions are made confidently while incorrect tokens are assigned smaller confidences.
Practical Computation of the Gatekeeper Loss. We evaluate Gatekeeper once per mini-batch within the standard training loop—no auxiliary passes or data-set re-shuffling are required. Given a mini-batch $B=\{(x_{i},y_{i})\}_{i=1}^{N}$, we perform a single forward pass through $\mathcal{M}_{S}$. This allows us to obtain (i) a logit vector $z_{i}=f_{\theta}(\mathbf{x}_{i})$, and (ii) predicted labels $\hat{y_{i}}=\arg\max_{c}z_{i,c}$. Two binary masks $m^{\text{corr}}_{i}=I\{y_{i}=\hat{y}_{i}\}$ and $m^{\text{incorr}}_{i}=\neg m^{\text{corr}}_{i}$ are computed on-the-fly. The hybrid loss is then assembled in a fully vectorized manner with components $\mathcal{L}_{\text{corr}}=\frac{1}{N}\sum_{i}m^{\text{corr}}_{i}\operatorname{CE}(p_{i},y_{i})$ and $\mathcal{L}_{\text{incorr}}=\frac{1}{N}\sum_{i}m^{\text{incorr}}_{i}\operatorname{KL}(p_{i}||\mathcal{U})$ where $p_{i}=\operatorname{softmax}(z_{i})$. Because both masks and losses are computed inside the same tensor graph, back-propagation incurs only the cost of $\mathcal{O}(N\times C)$ element-wise operations—identical to a vanilla cross-entropy step. This single-pass design keeps the computational overhead negligible while guaranteeing that in the full loss $\mathcal{L}=\alpha\mathcal{L}_{\text{corr}}+(1-\alpha)\mathcal{L}_{\text{incorr}}$ every sample contributes to either $\mathcal{L}_{\text{corr}}$ or $\mathcal{L}_{\text{incorr}}$ in the same optimization step.
Stage 2 (Inference): Confidence Computation & Thresholding. After fine-tuning $\mathcal{M}_{S}$ with Gatekeeper, we apply standard confidence- and entropy-based techniques for model uncertainty to obtain a deferral signal. We use the selective prediction framework to determine whether a query point $\mathbf{x}\in\mathbb{R}^{D}$ should be accepted by $\mathcal{M}_{S}$ or routed to $\mathcal{M}_{L}$. Selective prediction alters the model inference stage by introducing a deferral state through a gating mechanism (El-Yaniv and Wiener, 2010). At its core, this mechanism relies on a deferral function $g:\mathbb{R}^{D}\rightarrow\mathbb{R}$ which determines if $\mathcal{M}_{S}$ should output a prediction for a sample $\mathbf{x}$ or defer to $\mathcal{M}_{L}$. Given a targeted acceptance threshold $\tau$, the resulting predictive model can be summarized as:
$(\mathcal{M}_{S},\mathcal{M}_{L},g)(\mathbf{x})=\begin{cases}\mathcal{M}_{S}(\mathbf{x})&g(\mathbf{x})\geq\tau\\ \mathcal{M}_{L}(\mathbf{x})&\text{otherwise.}\end{cases}$ (4)
Classification Models (Max Softmax). Let $\mathcal{M}_{S}$ produce a categorical distribution $\{p(y=c\mid\mathbf{x})\}_{c=1}^{C}$ over $C$ classes. Then we define the gating function as
$g_{\text{CL}}(\mathbf{x})\;=\;\max_{1\,\leq\,c\,\leq\,C}\;p\bigl(y=c\,\big|\,\mathbf{x}\bigr).$ (5)
Token-based Models (Negative Predictive Entropy). Let $\mathcal{M}_{S}$ produce a sequence of categorical distributions $\{p(y_{t}=c\mid\mathbf{x})\}_{c=1}^{C}$ for each token index $t\in T$. Then we define the gating function as
$\footnotesize g_{\text{NENT}}(\mathbf{x})=\;\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{C}p\bigl(y_{t}=c\,\big|\,\mathbf{x}\bigr)\,\log p\bigl(y_{t}=c\,\big|\,\mathbf{x}\bigr),$ (6)
where $y_{t}\in[C]$ is the predicted token at time step $t$, $p(y_{t}=c\mid\mathbf{x})$ is the (conditional) probability of token $k$ at step $t$, and $T$ is the total number of token positions for the sequence. Across both model classes, higher values of $g_{\text{CL}}$ or $g_{\text{NENT}}$ indicate higher prediction confidence (i.e., lower uncertainty).
4 Experiments
In this section, we detail the experiments used to evaluate the effectiveness of Gatekeeper across three distinct model classes: encoder-only classification models, decoder-only language models, and encoder-decoder vision-language models. Each setup involves a cascade where a smaller model can defer inputs to a larger, more capable model. Our method and all competing baselines are applied on top of the same training/fine-tuning protocol used for obtaining the initial $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$ models.
4.1 Encoder-only Setup (Classification Models)
We comprehensively assess the performance of Gatekeeper across different model architectures and task types, starting with image classification. We train both a large model and a small model on the following datasets: CIFAR-10/100 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and TinyImageNet200 (Le and Yang, 2015). For both CIFAR datasets we use a ResNet-18 (He et al., 2016) as $\mathcal{M}_{L}$ and a custom CNN as $\mathcal{M}_{S}$. For Food-101 and TinyImageNet200 we instead use a ResNet-50 (He et al., 2016) as $\mathcal{M}_{L}$ and a Mobilenet V3 Small (Howard et al., 2019) as $\mathcal{M}_{S}$, where the latter is trained using knowledge distillation from the big model.
@@ 4.1 Encoder-only Setup (Classification Models)
Evaluation Metrics. We measure the performance of Gatekeeper and the resulting deferral function $g(\cdot)$ using the following performance metrics (see example in Figure 3 for an overview):
  1. 1.
    The Distributional Overlap of Confidences of Correct and Incorrect Predictions $s_{o}$ is defined as the integral of the minimum of the probability density functions (PDFs) of confidence scores for correctly classified samples, $\hat{p}_{\text{corr}}(c)$, and incorrectly classified samples, $\hat{p}_{\text{incorr}}(c)$ (see Figure 3a). Formally, given the confidence sets $\mathcal{C}_{\text{corr}}$ and $\mathcal{C}_{\text{incorr}}$, the overlap $s_{o}$ is computed as
    $s_{o}=\int_{0}^{1}\min\left\{\hat{p}_{\text{corr}}(c),\ \hat{p}_{\text{incorr}}(c)\right\}\,\mathrm{d}c,$ (7)
    where the PDFs are estimated using Kernel Density Estimation (KDE). If $s_{o}=1$, then $\mathcal{M}_{S}$ cannot distinguish the confidence distribution of correct and incorrect predictions; if $s_{o}=0$, then $\mathcal{M}_{S}$ can perfectly separate correct and incorrect predictions. Note that a related way of capturing the distributional separability is given by the Area Under the Receiver Operating Characteristic Curve (AUROC) which we discuss in Appendix C.3.
  2. 2.
    Deferral Performance $s_{d}$: To formally quantify how well $\mathcal{M}_{S}$ defers difficult inputs to $\mathcal{M}_{L}$, we examine the joint performance across all possible deferral ratios $r\in[0,1]$, where $r$ denotes the fraction of inputs sent to $\mathcal{M}_{L}$ based on a particular threshold $\tau$ (recall Equation (4)). Figure 3 b) illustrates how, as $r$ increases from 0 to $1$, the overall (joint) accuracy $\text{acc}(r)$ increases from the accuracy of $\mathcal{M}_{S}$ (yellow circle, no deferral) to the accuracy of $\mathcal{M}_{L}$ (purple square, full deferral). Useful deferral models are constrained to operate between random deferral ($\mathrm{acc}_{\mathrm{rand}}$, red dotted line) and ideal deferral ($\mathrm{acc}_{\mathrm{ideal}}$, green dashed line). The ideal deferral $\mathrm{acc}_{\mathrm{ideal}}$ corresponds to the oracle solution that perfectly defers examples misclassified by $\mathcal{M}_{S}$ and we discuss its exact functional form in Appendix B.3. We also define the realized deferral curve, $\mathrm{acc}_{\mathrm{real}}$, as the joint accuracy obtained under the learned deferral strategy $g(\cdot)$ employed by $\mathcal{M}_{S}$ and $\mathcal{M}_{L}$. The deferral performance metric $s_{d}$ is then given as:
    $\small s_{d}=\frac{A_{\mathrm{perf}}}{A_{\mathrm{useful}}}=\frac{\int_{0}^{1}\left(\mathrm{acc}_{\mathrm{real}}(r)-\mathrm{acc}_{\mathrm{rand}}(r)\right)\,\mathrm{d}r}{\int_{0}^{1}\left(\mathrm{acc}_{\mathrm{ideal}}(r)-\mathrm{acc}_{\mathrm{rand}}(r)\right)\,\mathrm{d}r}.$ (8)
    This ratio quantifies the fraction of the potential improvement over random deferral that has been realized by the achieved deferral strategy. Note that $s_{d}=1$ indicates perfect deferral, matching the ideal strategy, while an $s_{d}=0$ implies no improvement over random deferral.
  3. 3.
    Accuracy of the small model $\text{acc}(\mathcal{M}_{S})$: Finally, since Gatekeeper emphasizes patterns for distinguishing correct/incorrect examples, the model is no longer encouraged to minimize the classification loss over the full population. As a result, improving on the correct/incorrect separation task can lead to changes in utility over the full data distribution. Hence, practically useful deferral methods need to balance both deferral performance and the accuracy of $\mathcal{M}_{S}$.
[…]
This result highlights a critical trade-off which is directly controlled by $\alpha$:
How strongly do we want to degrade model performance over the full data distribution in order to obtain a better deferral model?
We note that this compromise between raw model utility and deferral performance is not surprising and similar trade-offs exist in fairness (Dutta et al., 2020; Yaghini et al., 2023) and privacy (Abadi et al., 2016; Rabanser et al., 2023). We study this trade-off explicitly in Figure 5 showing (i) a clear negative correlation between deferral performance and the small model’s accuracy; and (ii) a clear positive correlation between the overlap of correct/incorrect confidences and the accuracy of $\mathcal{M}_{S}$.
4.2 Decoder-Only Setup (LLMs)
In the decoder-only setup, we explore the application of LLMs. Our primary models of interest are the scalable LMs from the Gemma model class (GemmaTeam et al., 2024). We choose Gemma2B as $\mathcal{M}_{S}$ and Gemma7B as $\mathcal{M}_{L}$. Similar to the encoder-only setup, we employ smaller LMs as the initial classifiers to manage simpler next-token prediction tasks. The deferral strategy involves routing only those token sequences that exhibit high uncertainty—as determined by high predictive entropy—to the more powerful model $\mathcal{M}_{L}$.
@@ 4.2 Decoder-Only Setup (LLMs)
Figure 5: Performance trade-off between small model accuracy $\text{acc}(\mathcal{M}_{S})$ and deferral evaluation metrics. The baseline model obtained without fine-tuning using Gatekeeper is often the most accurate model over the full data distribution. With the introduction of Gatekeeper we can improve distinguishability of correct/incorrect predictions (top) as well as deferral performance (bottom) at the expense of model utility. Successful cascading solutions in practice need to balance both maintaining high model accuracy and deferral performance.
Our experiments begin by taking the instruction-tuned checkpoints of Gemma2B and Gemma7B and fine-tuning both models on the training split of each dataset to ensure that the model (i) performs well on the task and (ii) is familiar with the desired response format. This step is performed using standard supervised fine-tuning. Next, we fine-tune $\mathcal{M}_{S}$ with Gatekeeper on the same training split to reduce confidence on incorrect next-token predictions. Finally, we evaluate the model trained with Gatekeeper on a validation split. The datasets used are ARC-e/c (Clark et al., 2018), MMLU (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021). The evaluation metrics for our LLM experiments match those used in Section 4.1.
@@
4.3 Encoder-Decoder Setup (VLMs)
Finally, we examine models with both visual and textual processing capabilities, ideal for tasks requiring joint image understanding and language generation. We use the PaliGemma (Steiner et al., 2024) model family—encoder-decoder models designed for VL tasks such as image captioning, visual question answering, and descriptive image classification. The encoder processes input images into rich feature representations, while the decoder generates textual outputs. We use PaliGemma1B as $\mathcal{M}_{S}$ and PaliGemma7B as $\mathcal{M}_{L}$. Our deferral strategy runs the smaller VLM on all inputs and only defers to the more bigger 7B model when $\mathcal{M}_{S}$ ’s predictive entropy falls below a set threshold.
Similar to our experiments on LMs in Section 4.2, we employ two stages of fine-tuning. First, we take the instruction-tuned checkpoints of PaliGemma1B and PaliGemma7B and then fine-tune both models on the training split of a given dataset. Next, we fine-tune only $\mathcal{M}_{S}$ using Gatekeeper before evaluating the model on a validation/test split of the dataset. The datasets we consider are two classification datasets (VQAv2 (Goyal et al., 2017), AI2D (Hiippala et al., 2021)) and two captioning datasets (Cococap (Lin et al., 2014), Screen2Words (Wang et al., 2021)). This allows us to evaluate Gatekeeper in closed-form vision-language classification setups and open-form text generation.
@@ 4.3 Encoder-Decoder Setup (VLMs)
Factuality Scoring.
For classification tasks we apply our analysis in the same way as in Section 4.2. However, for captioning datasets we need to evaluate the quality of a caption generated by PaliGemma. To do that, we compute a factuality score which judges whether the generated caption is semantically coherent with respect to a reference caption using the Gemini LLM (GeminiTeam et al., 2023). Specifically, the Gemini LLM is prompted with an instruction of the form: “Are these captions semantically equivalent?”, followed by both the candidate caption and the reference caption. The model then responds with either “Yes” or “No”. Finally, we compute the log-likelihood of each response and normalize it to a probability, reflecting the LLM’s confidence in the captions being factually aligned. We detail this process in Appendix C.4 and denote the factuality score for input point $\mathbf{x}_{i}$ with candidate caption $\hat{\mathbf{y}}_{i}$ and ground truth caption $\mathbf{y}_{i}$ as $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$.
Measuring Correlation Between Factuality and Negative Predictive Entropy. Since the result of evaluating $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i})$ is no longer binary, our evaluation metrics which previously relied on accuracy cannot be used directly to evaluate deferral performance and the correct/incorrect entropy distribution separation. We address this issue by replacing the distributional overlap computation with the Pearson correlation $\rho(g_{\text{NENT}}(\mathbf{x}_{i}),s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i}))$ between the negative predictive entropy of a caption $g_{\text{NENT}}(\mathbf{x}_{i})$ and its associated factuality score $s_{\text{Fac}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{i}))$. We also adapt our deferral performance metric from Equation (8) to rely on factuality measures instead of accuracy.
[…]